OpenStudy (anonymous):

Can someone explain how to solve this equation??? Medals will be given log 2 + log(4x - 1) = 3

OpenStudy (ankitshaw):

i hope you know the properties of log .. log m + log n = logmn

OpenStudy (anonymous):

i kinda do

OpenStudy (ankitshaw):

use it u will get log (8x -2)= 3 sice base is 10 .. so. 10^3= 8x -2

OpenStudy (anonymous):

Can you explain how you got that answer?

OpenStudy (anonymous):

@ankitshaw

OpenStudy (ankitshaw):

log(8x-2) = \[\log_{10}(8x-2) \] we know that \[a^b = c \] can be written as \[\log_{a} c =b\]

OpenStudy (ankitshaw):

i hope now u will have no problem solving

OpenStudy (anonymous):

Thank you :)

OpenStudy (ankitshaw):

welcome